OpenStudy (anonymous):

lim h ->-2 (h^3 +8)/(h+2) can anyone help me?

jimthompson5910 (jim_thompson5910):

hint: a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (anonymous):

Jim, can you elaborate any?

jimthompson5910 (jim_thompson5910):

since 8 = 2^3 this means that for h^3 + 8 = h^3 + 2^3, we can see that a = h b = 2

jimthompson5910 (jim_thompson5910):

therefore, h^3 + 8 factors to (h+2)(h^2 - 2h + 4) using the first rule I posted

jimthompson5910 (jim_thompson5910):

so \[\large \frac{h^3 + 8}{h+2}\] becomes \[\large \frac{(h+2)(h^2 - 2h + 4)}{h+2}\]

OpenStudy (anonymous):

so then I just replace H with -2 and simplify it?

jimthompson5910 (jim_thompson5910):

simplify, then plug in h = -2

OpenStudy (anonymous):

umm, I get 0/0

OpenStudy (anonymous):

so is it DNE?

jimthompson5910 (jim_thompson5910):

look at \[\large \frac{(h+2)(h^2 - 2h + 4)}{h+2}\] notice anything cancelling?

OpenStudy (anonymous):

so it's h^2 - 2h + 4? so 4+4+4 so the limit is 12?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Wonderful! one more question

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

http://puu.sh/4InZm.jpg

OpenStudy (anonymous):

limit is -6 correct?

jimthompson5910 (jim_thompson5910):

hmm not sure about the double bar notation

jimthompson5910 (jim_thompson5910):

looks like absolute value, but why did they use 2 pairs instead of 1?

OpenStudy (anonymous):

greatest integer function, the thing when you round up

jimthompson5910 (jim_thompson5910):

oh ok

OpenStudy (anonymous):

like 2.01 = 3

jimthompson5910 (jim_thompson5910):

i gotcha

jimthompson5910 (jim_thompson5910):

aka, the ceiling function

jimthompson5910 (jim_thompson5910):

or wait, is that the floor, this page says otherwise http://www.mathwords.com/f/floor_function.htm

OpenStudy (anonymous):

it's the way that graph looks

OpenStudy (anonymous):

wait no it rounds up not down

jimthompson5910 (jim_thompson5910):

hmm odd

jimthompson5910 (jim_thompson5910):

that page says to round down

OpenStudy (anonymous):

I think that's a different version

jimthompson5910 (jim_thompson5910):

yeah it's a bit ambiguous

jimthompson5910 (jim_thompson5910):

I prefer floor and ceiling notations

jimthompson5910 (jim_thompson5910):

they are definitely more clear

myininaya (myininaya):

Are you doing problems from a book? If so, what is the way the function is defined there?

OpenStudy (anonymous):

also, is this: http://puu.sh/4IonL.png zero?

jimthompson5910 (jim_thompson5910):

0 is correct

OpenStudy (anonymous):

also, http://puu.sh/4Iovu.jpg is zero as well??

jimthompson5910 (jim_thompson5910):

correct again

OpenStudy (anonymous):

can you help me for this? http://puu.sh/4IoIm.jpg

jimthompson5910 (jim_thompson5910):

hint: rationalize the numerator by multiplying top and bottom by \(\large \sqrt{x}+5\) this is the conjugate of the numerator

OpenStudy (anonymous):

is the answer 0 again?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

what is the conjugate?

jimthompson5910 (jim_thompson5910):

in general if you have a+b the conjugate is a-b

jimthompson5910 (jim_thompson5910):

when you multiply the two, you get a difference of squares (a+b)(a-b) = a^2 - b^2

jimthompson5910 (jim_thompson5910):

so the conjugate of \(\large \sqrt{x}-5\) is \(\large \sqrt{x}+5\)

jimthompson5910 (jim_thompson5910):

multiply the two to get ????

OpenStudy (anonymous):

x - 25?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now multiply the denominator by \(\large \sqrt{x}+5\) this is so you keep things balanced

jimthompson5910 (jim_thompson5910):

notice how the x-25 terms will cancel

jimthompson5910 (jim_thompson5910):

this leaves you with \[\large \frac{1}{\sqrt{x}+5}\]

OpenStudy (anonymous):

is the limit 1/10?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

okay, so for this one http://puu.sh/4IpJO.jpg A and B are 0 and C is 2, correct?

jimthompson5910 (jim_thompson5910):

this will be the last one I help you with today

OpenStudy (anonymous):

that's fine :D

jimthompson5910 (jim_thompson5910):

all 3 answers you got are incorrect unfortunately

jimthompson5910 (jim_thompson5910):

try plugging in x = 3.01, then plug in x = 3.001

OpenStudy (anonymous):

the graph looks like http://puu.sh/4IpJO.jpg correct?

jimthompson5910 (jim_thompson5910):

what do you get as you get closer to x = 3?

jimthompson5910 (jim_thompson5910):

there is no graph in that post

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

http://puu.sh/4IpZQ.png

jimthompson5910 (jim_thompson5910):

much better

jimthompson5910 (jim_thompson5910):

so as you approach x = 3 from the right, where do you get close to (in terms of y)?

OpenStudy (anonymous):

shoot I made a silly mistake. A and B are 1?

jimthompson5910 (jim_thompson5910):

C is also 1 because the left and right hand limits are both 1

jimthompson5910 (jim_thompson5910):

remember that the limit only exists if the left and right limits are equal in this case, the limiting value equals the left and the right hand limits

OpenStudy (anonymous):

Wonderful! thank you so much for your help, I think I understand it better now

jimthompson5910 (jim_thompson5910):

that's great